# KTWIN: A Serverless Kubernetes-based Digital Twin Platform

### v1

@@ 5 Evaluation Methodology and Results
In this section, the evaluation methodology of this research is described. The evaluation involves designing, implementing, and deploying a Smart City Digital Twin use case using KTWIN. First, this section describes in more detail the ontology definitions for this specific use case, followed by the different evaluation scenarios, and the evaluation setup. The final sections present the evaluation results of the executed experiments, including service scalability and the KTWIN serverless capabilities analysis.
@@ 5.1 Smart City Evaluation Use Case
In order to evaluate the proposed solution, a Smart City use case was designed, implemented, and deployed using KTWIN. The selected use case was based on the DTDL ontology for Smart Cities developed by Open Agile Smart Cities (OASC) and Microsoft [32]. The open-source ontology model was publicly released to accelerate the development of Digital Twin-based solutions for Smart Cities. The DTDL is composed of ontology-based definitions from NGSI-LD information model specification [33] and ETSI SAREF ontologies for Smart Cities (Saref4City) [34]. The final model comprises entities in Urban Mobility, Environment, Waste, Parking, Buildings, Parks, Ports, City Objects, Administrative Areas, and more. Figure 5 shows the Smart City DTDL ontology.
@@
The open source DTDL ontology was converted to KTWIN YAML definitions files using KTWIN CLI implemented as part of the prototype. The YAML files mostly include the entity’s data model definitions, such as attributes, data-type definitions, and existing interface relationships. The Smart City design effort also included the creation of additional properties and commands, not available in the original DTDL, to fulfill the events definition scenarios described in the following sections. In addition, the conversion effort included the creation of Twin Instances definition files, not available in the original DTDL specification. Finally, a set of container-based applications were developed using KTWIN SDK, presented in the next section, as stateless functions to handle events for the deployed Twin Instances, and defined in the deployment spec files. The same artifacts included parameterized settings to configure auto-scaling policies and control the allocated CPU and memory. The generated artifacts for the prototype evaluation are available on GitHub  [^3] https://github.com/Open-Digital-Twin/ktwin-article.
Figure 5: Smart City DTDL Ontology Diagram.
@@ 5.2 Evaluation Scenarios and Metrics
The evaluation scenario of this research aims to reproduce a set of the rich number of events that happen during a day-in-the-life of a Smart City. The scenarios were built based on the DTDL presented in the previous section. Each DTDL entity was mapped to a Twin Interface and a set of Twin Instances using the KTWIN CRDs required by the KTWIN Operator. The evaluation scenario reproduces the events that occur within 24 hours of activities in a modern city and considers the different frequencies of events depending on their nature. A summary of the corresponding Twin Interfaces and the nature of their events is presented in more detail as follows.
  • 1.
    City: represents the twin graph’s higher-level entity, the city itself. A city is a large human settlement. A city contains administrative regions such as neighborhoods or districts.
  • 2.
    Neighborhood: A neighborhood is a geographically localized community within a larger city, town, suburb, or rural area. The neighborhood can compute and aggregate data from other instances, such as the air quality level of a certain region or the number of available public parking spots in a determined place in the city.
  • 3.
    Smart Pole: A smart pole is localized in a determined street that belongs to a neighborhood. The pole contains a set of sensors that allow monitoring of the environmental quality variables in the area, such as air quality, noise level, crowd and traffic flow, or weather measurements. A Smart Pole can also be connected to an Electric Vehicle (EV) Charging Station.
  • 4.
    Air Quality Observed: It represents the air quality observed in a certain region of the city in the determined period. The air quality observed registers the density of different gases in a certain area such as Carbon dioxide (CO2), Carbon monoxide (CO), and Sulfur dioxide (SO2). The implemented service classifies the Air Quality Index (AQI) [35] of the area into Good, Moderate, Unhealthy for Sensitive Groups, Unhealthy, Very Unhealthy, and Hazardous based on the gas density observed by the sensors. The service also updates the air quality index of the neighborhood interface to which it is related.
  • 5.
    Weather Observed: The weather variables observations of a smart pole region are registered in the Weather Observed interface. The interface contains weather-related information such as temperature, precipitation, humidity, snow height, wind direction and speed, and atmospheric pressure. The implemented service aggregates the measured data and computes additional information such as the dew point, feel likes temperature, and pressure tendency.
  • 6.
    Crowd Flow Observed: The interface is used to monitor the crowd congestion in the city during the day. It can be used to monitor the number of people, the flow direction, and the average crowd speed observed in a certain subway station or bus terminal, for example.
  • 7.
    Traffic Flow Observed: The Traffic Flow information is an observation of traffic conditions at a certain place and time. The entity contains information such as the distance and the average headway time between two consecutive vehicles to determine the region’s congestion level.
  • 8.
    Noise Level Observed: The noise level observed entity contains information about the sound pressure level in a certain region of the city.
  • 9.
    EV Charging Station: This entity represents a smart station connected to a Smart Pole that allows electric vehicles to charge their batteries at any time in the day. A charging station may accept only a set of vehicle types. Additionally, it has its own working status such as full usage, not in use, or out of service.
  • 10.
    Streetlights: Streetlights are associated with the city pole and monitor the lighting level in different regions of the city. In case there is a defective lamp, this information can be used to replace it.
  • 11.
    Parking Spot: A Parking Spot represents a place for parking vehicles within the city. The spot can have its status change over the day from free, occupied or closed. A Parking Spot can belong to an Off-Street Parking or On Street Parking. Parking Groups can be used optionally to group parking spots by different criteria such as the floor number, in case of private building park, or street, in case of public parking.
  • 12.
    Device: A Device can be any kind of sensor, actuator, or camera that generates data for the system. Devices are associated with the majority of the entities previously listed. For instance, a device was involved in order to measure the air quality of a certain area of the city or to report the parking slot availability. The device interface is designed to monitor different operational data from sensors, for example, each device contains its own battery level that can be measured to monitor the device’s health.
Figure 6 shows a code example of the Air Quality Observed function implemented using the KTWIN SDK in Golang with some redacted lines. The container-based application exposes an HTTP endpoint and implements a HandleEvent method that converts the HTTP request into a Twin Event structure compatible with Cloud Event spec. The function calculates the Air Quality Index of the corresponding area and publishes a command to the smart pole instance using the Twin Graph loaded from Graph Cache.
func HandleEvent(event *ktwin.TwinEvent) error {
return kevent.HandleEvent(event, TWIN_INTERFACE_AIR_QUALITY_OBSERVED, handleAirQualityObservedEvent)
func handleAirQualityObservedEvent(event *ktwin.TwinEvent) error {
var airQualityObserved model.AirQualityEvent
err := event.ToModel(&airQualityObserved)
airQualityObserved.CalculateAqiLevel()
event.SetData(airQualityObserved)
err = keventstore.UpdateTwinEvent(event)
allLevels := []model.AQICategory{airQualityObserved.COAqiLevel,
airQualityObserved.SO2AqiLevel, airQualityObserved.O3AqiLevel}
var command model.UpdateAirQualityIndexCommand
updateAirQualityIndexCommand.SetAqiLevel(allLevels)
err = kcommand.PublishCommand(TWIN_COMMAND_AIR_QUALITY_CITY_POLE_UPDATE_AIR_QUALITY_INDEX, command,
TWIN_COMMAND_AIR_QUALITY_CITY_POLE_RELATIONSHIP_NAME, event.TwinInstance, *twinGraph)
Figure 6: Air Quality Observed function implemented using Golang SDK.
The simulation scenario was designed to reproduce the 24-hour daily routine of a Smart City. The simulation design took into account (1) events that are generated throughout the day on a defined frequency, contributing to a fixed data traffic, and (2) events that are generated at a specific moment of the day and do not follow a well-established frequency or are generated because some event criteria in the system were reached, contributing to unpredictable data traffic.
@@
In this context, the environment-measured data such as air quality index, city weather, temperature, and noise level observed are events generated at a fixed frequency throughout the full day. However, events directly related to city mobility, such as observed crowds and traffic flows, changes in parking spot availability, and the monitoring of available electric vehicle charging stations, present variable frequencies during the day depending on the city’s peak and off-peak hours. Another example of events that present unpredictable traffic is the monitoring of the battery level of sensor devices - each device has its battery level consumption and recharge particularities, so it is difficult to predict when this kind of event will be generated. In-system events, such as the Air Quality Index warning alert when some Smart Pole identifies a low-quality air measurement, are also part of the evaluation scenario.
@@
Table 3 shows the events implemented and analyzed in the evaluation scenario with their corresponding time interval variation for peak and off-peak hours. The selected event interval considered for simulation compression was from 24 hours to 24 minutes, which means that a 10-second interval represents ten minutes in the real world, allowing the execution of several evaluation experiments in a reduced timeframe.
@@
The experiment divides the day simulation into six windows of four hours each, corresponding to 240 seconds in the simulation time. Each event type may differ in frequency within each time window, allowing the reproduction of the variation of events at different moments of the day, such as a city’s peak and off-peak hours. In addition to the generated MQTT device events, Twin Services generates virtual events back to the edge devices and publishes events for command executions and data persistence.
@@
The execution of the experiment was divided into two steps. First, the above scenarios were executed 3 times each with autoscale and scale-to-zero features enabled to verify how the prototype scales for different city sizes using Knative Serverless functions. The minimum number of pods per Twin Instances was set to zero and the maximum was set to 18. The Event Store service thresholds were set to 1 and 25, because it has a higher throughput. The selected upper thresholds correspond to the maximum amount of pods that could be created considering the memory limits of the nodes set to each they were deployed. The autoscale settings were configured using the concurrency metric, which determines the number of simultaneous requests that each replica of an application can process at any given time, as the value 5. Later, the scenarios of 1 and 5 neighborhoods were executed 3 times each with a fixed number of pods to reproduce scenarios of over-provisioning computing resources (160 pods, 14 per Twin Instance + 20 for the Event Store) and under-provisioning computing resources (13 pods, 1 per Twin Instance + 3 for the Event Store). Each simulation aimed to reproduce 24 hours of Smart City events in a 24-minute timeframe. The evaluation metrics collected in the experiments include memory usage, CPU usage, and response time.
@@ 5.4.1 Evaluation Scenarios Overview
In the 1-neighborhood scenario, the range of processed events ranges from 13 events per second to 101 events per second, with a median of 67 events per second. In the 5-neighborhood simulation, the median is approximately 5 times the previous scenario, 334 events per second, while the range is between 107 and 475. For the remaining scenarios, the median is 666 and 1301 events per second, while the ranges are between 168 and 1062, and 494 and 1725 events per second. Not all events processed by the broker have devices as sources. In-cluster events to propagate commands, perform data aggregation, and events propagated to the event store are also accounted for in these experiments.
@@
RabbitMQ performance is related to the number of queues and bindings created within the broker. Because KTWIN handles routing rules on the Twin Interface level, the number of created queues is proportional to the number of defined interfaces. Figure 8(b) shows the number of RabbitMQ resources for all experiments. Independently of the experiment sizing, the number of RabbitMQ resources remains constant, reducing the broker memory footprint and the Kubernetes resource storage: the RabbitMQ internal resources are instantiated using the Kubernetes RabbitMQ Topology Operator. During some early stages of the prototype evaluation, we verified a higher memory consumption in RabbitMQ when implementing the routing rules on top of Twin Instances. In the latest KTWIN prototype, the RabbitMQ instance contains two exchanges used by the broker ingress component in which Twin Services publish all their events and the standard RabbitMQ MQTT exchange to which devices publish MQTT messages. In total, the simulation events were routed through 13 different queues.
@@ 5.4.2 Services Scalability Analysis
(a) Neighborhood-scenario
(b) Neighborhoods-scenario
(c) 0 Neighborhoods-scenario
(d) 0 Neighborhoods-scenario
@@ 5.4.3 Serverless Capabilities Analysis
KTWIN Serverless capabilities provide several benefits to end users. The platform provides operational complexity and cost reduction allowing domain experts to define their own ontology-based Digital Twin definitions and perform rapid development of Twin Interface functions without requiring deep knowledge of infrastructure management. KTWIN Serverless capabilities also include the possibility to offer the solution as a Pay-as-you-Go subscription model in which users only pay for the compute time, the number of events processed, the twin size and the amount of data storage. In addition, KTWIN provides parameterized scalability of components with automatic scaling based on demand and the dynamic allocation of resources to ensure optimal performance during varying loads. These aspects enable KTWIN to be the foundation of a Serverless Platform for Digital Twins. In order to evaluate the functions’ scalability and the achievable cost savings, we compare the 1-neighborhood and 5-neighborhood experiments against scenarios with fixed amounts of under-provisioned resources (the number of pods per Twin Interface was set to 1), and over-provisioned scenarios (14 pods per Twin Interface).
@@
Figure 16(b) shows that the CPU used by the services for the scenarios with automatic scale remains under 20% of usage, almost the same range registered in the overprovisioned scenario. In contrast, CPU usage in the under-provisioned 5-neighborhood scenario stays above 50% during most of the experiment period. The chart presented in 16(b) suggests that auto-scaler scenarios can allocate less resource usage and use the allocated ones more efficiently.
@@
(b) Services CPU used/requested.
@@
The same resource usage efficiency can be verified in the memory usage charts. Figure 17 shows that the amount of allocated memory remains far below the scenarios of over-provisioned resources, while the percentage of usage memory presents a better efficiency compared to the under-provisioned scenario. The allocated memory savings range between 64% and 89% compared to the over-provisioned configuration.
@@
The results show that KTWIN achieves resource usage efficiency closer to the under-provisioned scenario, while still providing performance near the over-provisioned scenario. In contrast, the scenarios with autoscale and scale-to-zero face the cold start problem, resulting in some initial latency before resolving bursts of unpredicted events. In this context, high-priority events that require to be processed within an established timeframe should not consider scale-to-zero as an option.
(a) Services memory requested.
(b) Services memory used/requested.
Figure 17: Memory requested and used by services.

### v3

@@ 6 Evaluation Methodology and Results
In this section, the evaluation methodology of this research is described. The evaluation involves designing, implementing, and deploying a Smart City Digital Twin use case using KTWIN. First, this section describes in more detail the ontology of the selected use case as well as the implementation details, followed by the different evaluation scenarios, and the evaluation setup. The final sections present the evaluation results of the executed experiments, including service scalability and the KTWIN Serverless capabilities analysis.
@@ 6.1 Smart City Evaluation Use Case
In order to evaluate the proposed solution, a Smart City use case was designed, implemented, and deployed using KTWIN. The selected use case was based on the DTDL ontology for Smart Cities developed by Open Agile Smart Cities (OASC) and Microsoft [38]. The open-source ontology model was publicly released to accelerate the development of Digital Twin-based solutions for Smart Cities. The DTDL is composed of ontology-based definitions from NGSI-LD information model specification [39] and ETSI SAREF ontologies for Smart Cities (Saref4City) [40]. The final model comprises entities in Urban Mobility, Environment, Waste, Parking, Buildings, Parks, Ports, City Objects, Administrative Areas, and more. A summary of the corresponding group of Twin Interfaces created for the evaluation scenarios is presented in more detail as follows.
Administrative Area: set of entities that represent the organizational structure of a city. At the highest level, the City represents a large human settlement comprising administrative regions such as neighborhoods or districts. A Neighborhood is a geographically localized community within a city, suburb, or rural area, functioning as a smaller administrative unit. Neighborhoods can compute and aggregate data from various entities, such as monitoring air quality in specific regions or tracking the availability of public parking spots in designated areas.
Utilities: this category encompasses a set of entities designed to provide essential resources for residents. A Smart Pole is located on a specific street within a neighborhood and is equipped with various sensors to monitor environmental quality metrics, including air quality, noise levels, crowd density, traffic flow, and weather conditions. Additionally, a Smart Pole can integrate with an Electric Vehicle (EV) Charging Station to support sustainable transportation. Streetlights are associated with Smart Poles, enabling the monitoring of lighting levels across different city regions. If a lamp becomes defective, the system can detect it and provide actionable information for timely replacement.
Environment: set of entities used to measure the environment quality observed in certain neighborhood of the city. The Air Quality Observed represents the air quality in a certain region of the city in the determined period. It registers the density of different gases in a certain area such as Carbon dioxide (CO2), Carbon monoxide (CO), and Sulfur dioxide (SO2). The implemented service classifies the Air Quality Index (AQI) [41] of the area into Good, Moderate, Unhealthy for Sensitive Groups, Unhealthy, Very Unhealthy, and Hazardous based on the gas density observed by the sensors. The service also updates the air quality index of the neighborhood interface to which it is related. The weather variables observations of a smart pole region are registered in the Weather Observed interface. The interface contains weather-related information such as temperature, precipitation, humidity, snow height, wind direction and speed, and atmospheric pressure. The implemented service aggregates the measured data and computes additional information such as the dew point, feel likes temperature, and pressure tendency. Lastly, the noise level observed entity contains information about the sound pressure level in a certain region of the city.
Mobility: set of entities used for monitoring and managing urban mobility. The Crowd Flow Observed entity tracks crowd congestion, including the number of people, flow direction, and average speed at locations like subway stations or bus terminals. Traffic Flow Observed monitors traffic conditions, providing data such as vehicle spacing and headway times to assess congestion levels. EV Charging Stations are smart facilities linked to Smart Poles, enabling electric vehicles to charge with status updates like available, in use, or out of service, and compatibility with specific vehicle types. Parking Spots represent designated parking areas, with dynamic statuses (free, occupied, or closed) and can be categorized as Off-Street (e.g., garages) or On-Street (e.g., public roads). Parking Groups optionally organize spots by criteria like floor or street.
Cross-Domain: a set of generic entities that can be shared across different domains. A Device can be any kind of sensor, actuator, or camera that generates data for the system. Devices are associated with the majority of the entities previously listed. For instance, a device was involved in order to measure the air quality of a certain area of the city or to report the parking slot availability. The device interface is designed to monitor different operational data from sensors, for example, each device contains its own battery level that can be measured to monitor the device’s health.
@@
The open source DTDL ontology was converted to KTWIN CRDs files using KTWIN CLI implemented as part of the prototype. Each DTDL entity was mapped to a Twin Interface and a set of Twin Instances, and later applied to the cluster to be provisioned by the KTWIN Operator. The CRDs files mostly include the entity’s data model definitions, such as attributes, data-type definitions, and existing interface relationships. The Smart City design effort also included the creation of additional properties and commands, not available in the original DTDL, to fulfill the events definition scenarios described in Table 3. Finally, a set of container-based applications were developed using KTWIN SDK, a library used to abstract event handling, propagation and persistence within KTWIN. The containers were implemented as stateless functions to handle events for the deployed Twin Instances, and defined in the deployment specification files. The same artifacts included parameterized settings to configure auto-scaling policies and control the allocated CPU and memory. The generated artifacts for the prototype evaluation are available on GitHub  [^3] https://github.com/Open-Digital-Twin/ktwin-article.
@@ 6.2 Evaluation Scenarios and Metrics
The evaluation scenario of this research aims to reproduce a set of the rich number of events that happen during a day-in-the-life of a Smart City. The simulation design took into account (1) events that are generated throughout the day on a defined frequency, contributing to a fixed data traffic, and (2) events that are generated at a specific moment of the day and do not follow a well-established frequency or are generated because some event criteria in the system were reached, contributing to unpredictable data traffic. In this context, the environment-measured data such as air quality index, city weather, temperature, and noise level observed are events generated at a fixed frequency throughout the full day. However, events directly related to city mobility, such as observed crowds and traffic flows, changes in parking spot availability, and the monitoring of available electric vehicle charging stations, present variable frequencies during the day depending on the city’s peak and off-peak hours. Another example of events that present unpredictable traffic is the monitoring of the battery level of sensor devices - each device has its battery level consumption and recharge particularities, so it is difficult to predict when this kind of event will be generated. In-system events, such as the Air Quality Index warning alert when some Smart Pole identifies a low-quality air measurement, are also part of the evaluation scenario.
@@
Table 3 shows the events implemented and analyzed in the evaluation scenario with their corresponding time interval variation for peak and off-peak hours. The selected event interval considered for simulation compression was from 24 hours to 24 minutes, which means that a 10-second interval represents ten minutes in the real world, allowing the execution of several evaluation experiments in a reduced timeframe. The experiment divides the day simulation into six windows of four hours each, corresponding to 240 seconds in the simulation time. Each event type may differ in frequency within each time window, allowing the reproduction of the variation of events at different moments of the day, such as a city’s peak and off-peak hours. In addition to the generated MQTT device events, Twin Services generates virtual events back to the edge devices and publishes events for command executions and data persistence.
@@
The execution of the experiment was divided into two steps. First, the above scenarios were executed 3 times each with autoscale and scale-to-zero features enabled to verify how the prototype scales for different city sizes using Knative Serverless functions. The minimum number of pods per Twin Instances was set to zero and the maximum was set to 18. The Event Store service thresholds were set to 1 and 25, because it has a higher throughput. The selected upper thresholds correspond to the maximum amount of pods that could be created considering the memory limits of the nodes set to each they were deployed. The autoscale settings were configured using the concurrency metric, which determines the number of simultaneous requests that each replica of an application can process at any given time, as the value 5. Later, the scenarios of 1 and 5 neighborhoods were executed 3 times each with a fixed number of pods to reproduce scenarios of over-provisioning computing resources (160 pods, 14 per Twin Instance + 20 for the Event Store) and under-provisioning computing resources (13 pods, 1 per Twin Instance + 3 for the Event Store). Each simulation aimed to reproduce 24 hours of Smart City events in a 24-minute timeframe.
The experiments focused on evaluating performance and resource utilization using a set of carefully selected metrics: the number of pods, CPU, memory, and response times. The number of pods metric represents the number of deployed instances in the system and is used to understand how the system scales service instances for different workload sizes. The CPU metrics include CPU requested, reflecting the amount of CPU requested for each service instance, and CPU usage, indicating the actual percentage of resource utilization. Similarly, memory metrics comprised memory requested, representing planned memory allocation, and memory usage, showing real consumption levels. To measure performance under varying loads, response time metrics were analyzed across multiple percentiles: p50 (median) to gauge typical user experience, p90 and p95 to capture performance under higher loads, and p99 to evaluate system behavior during peak or extreme conditions. Together, these metrics provided a comprehensive understanding of the system’s scalability, efficiency, and performance across deployment configurations and neighborhood sizes.
@@ 6.4.1 Evaluation Scenarios Overview
In the 1-neighborhood scenario, the number of processed events ranges from 13 events per second to 101 events per second, with a median of 67 events per second. In the 5-neighborhood simulation, the median is approximately 5 times the previous scenario, 334 events per second, while the range is between 107 and 475. For the remaining scenarios, the median is 666 and 1301 events per second, while the ranges are between 168 and 1062, and 494 and 1725 events per second. Not all events processed by the broker have devices as sources. In-cluster events to propagate commands, perform data aggregation, and events propagated to the event store are also accounted for in these experiments.
@@
RabbitMQ performance depends on the number of queues and bindings within the broker. By managing routing rules at the Twin Interface level, KTWIN ensures that the number of queues scales proportionally to the defined interfaces, demonstrating robust scalability across varying experiment sizes. Figure 6(b) illustrates the number of RabbitMQ resources used in all experiments. Regardless of experiment size, the number of RabbitMQ resources remains consistent, minimizing the broker’s memory footprint and reducing Kubernetes resource storage. In earlier prototype evaluations, several routing strategies were evaluated, such as the implementation of routing rules on top of Twin Instances, leading to increased RabbitMQ memory usage. In the latest KTWIN prototype, the broker instance is optimized to include only two exchanges: one for the broker ingress component, where all Cloud Events are published, and the MQTT exchange, where devices publish and subscribe to messages. The simulation events were effectively routed through only 13 queues regardless of the city size, showcasing KTWIN’s scalable and efficient routing capabilities.
@@ 6.4.2 Services Scalability Analysis
(a) 1 Neighborhood-scenario
(b) 5 Neighborhoods-scenario
(c) 10 Neighborhoods-scenario
(d) 20 Neighborhoods-scenario
@@ 6.4.3 Serverless Capabilities Analysis
KTWIN Serverless capabilities provide several benefits to end users. The platform provides operational complexity and cost reduction allowing domain experts to define their own ontology-based Digital Twin definitions and perform rapid development of Twin Interface functions without requiring deep knowledge of infrastructure management. KTWIN Serverless capabilities also include the possibility to offer the solution as a Pay-as-you-Go subscription model in which users only pay for the compute time, the number of events processed, the twin size and the amount of data storage. In addition, KTWIN provides parameterized scalability of components with automatic scaling based on demand and the dynamic allocation of resources to ensure optimal performance during varying loads. These aspects enable KTWIN to be the foundation of a Serverless Platform for Digital Twins. In order to evaluate the functions’ scalability and the achievable cost savings, we compare the 1-neighborhood and 5-neighborhoods experiments against scenarios with fixed amounts of under-provisioned resources (the number of pods per Twin Interface was set to 1), and over-provisioned scenarios (14 pods per Twin Interface).
@@
(b) Services memory requested.
(c) Services CPU used/requested.
(d) Services memory used/requested.
@@
Figure 14(c) shows that the CPU used by the services for the scenarios with automatic scale remains under 20% of usage, almost the same range registered in the overprovisioned scenario. In contrast, CPU usage in the under-provisioned 5-neighborhood scenario stays above 50% during most of the experiment period. The chart presented in 14(c) suggests that auto-scaler scenarios can allocate less resource usage and use the allocated ones more efficiently.
The same resource usage efficiency can be verified in the memory usage charts. Figures 14(b) and 14(d) show that the amount of allocated memory remains far below the scenarios of over-provisioned resources, while the percentage of usage memory presents a better efficiency compared to the under-provisioned scenario. The allocated memory savings range between 64% and 89% compared to the over-provisioned configuration.
@@
The results show that KTWIN achieves resource usage efficiency closer to the under-provisioned scenario, while still providing performance near the over-provisioned scenario. In contrast, the scenarios with autoscale and scale-to-zero face the cold start problem, resulting in some initial latency before resolving bursts of unpredicted events. In this context, high-priority events that require to be processed within an established timeframe should not consider scale-to-zero as an option. Since the experiments compared Serverless scenarios with under and over-provisioned configurations, the findings are not limited to the choice of Knative as the Serverless platform. Similar results are expected regardless of the selected platform.